# inkscapelatex=false

### v1

@@ Abstract.
A key component of transformers is the attention mechanism orchestrating how each token influences the propagation of every other token through a transformer. In this paper we provide a rigorous, mathematical analysis of the asymptotic properties of attention in transformers. Although we present several results based on different assumptions, all of them point to the same conclusion, all tokens asymptotically converge to each other, a phenomenon that has been empirically reported in the literature. Our findings are carefully compared with existing theoretical results and illustrated by simulations and experimental studies using the GPT-2 model.
@@ 1. Introduction
The incorporation of attention [1] in natural language processing was a significant breakthrough, particularly in the context of sequence-to-sequence models, enabling the creation of transformers [2] which revolutionized the field. Even initial transformer models such as GPT [3] or Bert [4] showed drastic improvements over previous approaches such as the Long Short-Term Memory model [5].
As practical applications of deep neural networks, such as image recognition [6], natural language processing [7], and autonomous driving [8], continue to advance, our understanding of these networks is struggling to keep pace [9]. This underscores the critical importance of our study, which aims to delve deeper into transformers and their dynamics. Our understanding of transformers is currently limited by their inherent complexity, making it challenging to comprehensively explain their behavior [10]. However, recent studies have shown the emergence of clusters of tokens empirically and theoretically [11; 12; 13; 14]. These findings suggest that without proper care, large transformers may collapse, a phenomenon where the tokens cluster, limiting the model’s ability to produce different outputs.
Our work was motivated by the paper [15] where a mathematical model for attention was proposed, based on prior work on similar models [16; 17], and investigated. The authors share the vision outlined in [15], a better understanding of the role and importance of attention mechanisms can be achieved through the study of mathematical models. Our contribution lies in bringing ideas developed by the control community, where the study of asymptotic properties of dynamical and control systems is a central preoccupation, to bear on this problem. While deferring to the next section a more detailed comparison between our results and those available in the literature, we emphasize here that, in contrast with [15; 13; 14], we do not rely on stochastic and/or mean-field techniques and rather adopt a geometric perspective drawing from control theory, e.g., from consensus dynamics on manifolds [18] such as spheres  [19; 20] and Input-to-State Stability [21; 22; 23].
Contributions and plan of the paper
The main contribution of this work is to provide a number of results, for a differential equation model of attention, showing that all tokens converge to a single cluster thereby leading to a collapse of the model. We use the term consensus equilibria to refer to such clusters as is done in the consensus literature [24; 25]. These results hold under different assumptions on the parameters of the model —namely, the query ($Q$), key ($K$) and value matrices ($U$), as well as the number of heads ($h$)— that are summarized in Table 1. More specifically, the paper is organized as follows:
  • In Section 2 we introduce the differential equation model for attention studied in this paper. Since layer normalization is part of the model, tokens will evolve on ellipsoids. As we are mainly concerned with attention, the model does not describe the effect of feedforward layers [^1] By feedforward layer we mean the layer implementing a perceptron, i.e., an elementwise activation function acting on an affine function of the input, see, e.g., [26]. in a transformer. Yet, we briefly discuss how the model can be extended to accommodate feedforward layers and the challenges brought by such extension.
  • Section 3 is devoted to the single-head case with $U$ being the identity and $P=Q^{\top}\,K$ being time invariant, positive definite, and symmetric. With Theorem 3.2, we prove that the dynamics of the transformer is a Riemannian gradient vector field, from which we conclude convergence to an equilibrium point (guaranteed to be of consensus type when $P$ is the identity) for every initial position of the tokens. Although the gradient nature of the dynamics, in this case, was already observed and exploited in [15], for the benefit of the readers we provide a formal proof of this fact in a slightly more general setting.
  • In Section 4 we show that tokens converge to a consensus equilibrium whenever their starting positions lie in the interior of some hemisphere of the ellipsoid. This is stated in Theorem 4.1, which holds for any number of heads and time varying matrix $P=Q^{\top}\,K$ provided that $U$ is the identity and $P$ is bounded and uniformly continuous as a function of the time. A similar result is reported in [15] under Lemma 4.2. However, its conclusions hold under the stronger assumptions that both $U$ and $P=Q^{\top}K$ are the identity matrix and there is a single attention head.
The previous results hold under no assumptions on the attention matrix other those induced by the assumptions on $P$. In the next sections, we focus on the auto-regressive case, also known as causal attention, where the self-attention matrix is lower triangular.
  • For the auto-regressive case with $U$ being the identity, the first token is fixed. In Section 5, we show that all tokens converge to the position of the first one for almost every initial position of the tokens. In fact, Theorem 5.1 ensures asymptotic stability of this consensus equilibrium. This holds for any number of heads and any time varying $P$ matrix provided it is bounded. Similar conclusions are reported under Theorem 4.1 in [27] by imposing stronger assumptions: time invariance of $P=Q^{\top}K$ and existence of a single attention head.
  • To conclude the theoretical part, Section 6 extends the previous result to the case where $U$ is a time invariant symmetric matrix and the multiplicity of its largest eigenvalue is one. Therefore, the corresponding eigenspace divides the sphere in two different hemispheres. Theorem 6.1 ensures that the tokens will converge to a consensus equilibrium (moreover, that equilibrium is asymptotically stable) if all the tokens start in one of those hemispheres. We were only able to establish this result for the single-head case although we believe it holds in greater generality. To the best of the author’s knowledge there is no result available in the literature for the case where $U$ is not the identity matrix although this is conjectured, but not proved, in [27].
  • In Section 7 we illustrate the theoretical results through simulations of the mathematical model for attention. We do this using a small number of tokens in low dimensions, for better visualization, as well as a number of tokens and dimension comparable to what is used in the GPT-2 model. We also report on several experiments with the GPT-2 model suggesting convergence to a consensus equilibria in more general situations than those captured by our theoretical results thus providing additional confirmation for model collapse.
Full attention Causal attention (auto-regressive)
Section §3 §4 §5 §6
# of heads $h=1$ $h\geq 1$ $h\geq 1$ $h=1$
$P(t)=Q(t)^{\top}\,K(t)$ Time invariant, symmetric, positive definite Time varying, uniformly continuous, bounded Time varying, bounded Time varying, bounded
$U(t)$ Identity Identity Identity Time invariant, symmetric
Result Theorem 3.2 Theorem 4.1 Theorem 5.1 Theorem 6.1
Statement Gradient flow, convergence to equilibrium Convergence to consensus Asympt. stability of consensus (determined by the first token) Asympt. stability of consensus (determined by eigenspace of largest eigenvalue of $U$)
Domain of attraction Whole sphere Some hemisphere Conull (complement of zero measure) Fixed hemisphere
@@
Table 1. Summary of the results presented in this work for several particular cases of the continuous model (7), where $Q(t)$, $K(t)$, and $U(t)$ denote the query, key and value matrices, respectively.
Notations
We use the letters $n,\ell,r$ and $s$ to denote natural numbers, i.e., elements of =$\mathbb{N}=\{1,2,\ldots\}$. The space of $r\times s$ real matrices is denoted by $\mathcal{M}_{r\times s}(\mathbb{R})$. In particular, $\mathbb{I}_{r}\in\mathcal{M}_{r\times r}(\mathbb{R})$ denotes the identity matrix. The elements of $\mathbb{R}^{n+1}$, denoted by ${\rm x}$, are regarded as column matrices, i.e., ${\rm x}\in\mathbb{R}^{n+1}\equiv\mathcal{M}_{(n+1)\times 1}(\mathbb{R})$. Tuples of $\ell$ elements of $\mathbb{R}^{n+1}$ are denoted by $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})$ (note the different font). When it is convenient, they will be regarded either as matrices $\mathbf{x}\in\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})$ or column matrices, i.e., $\mathbf{x}\in\mathcal{M}_{(n+1)\ell\times 1}(\mathbb{R})$.
The tangent space of a smooth manifold $M$ at $p\in M$ and its elements are denoted by $T_{p}M$ and $X_{p}\in T_{p}M$, respectively. The corresponding tangent bundle and the space of vector fields are denoted by $\pi_{M}:TM\to M$ and $\mathfrak{X}(M)=\Gamma(\pi_{M})$, respectively. In the same vein, the space of $k$-forms is denoted by $\Omega^{k}(M)$. Given another smooth manifold $N$ and a smooth map $\phi:M\to N$, i.e., $\phi\in C^{\infty}(M,N)$, its tangent map is denoted by $T\phi:TM\to TN$ while its pullback is denoted by $\phi^{*}:T^{*}N\to T^{*}M$.
@@
The inner product between the vectors $X_{p}\in T_{p}M$ and $Y_{p}\in T_{p}M$, according to a Riemannian metric $g$ on $M$, is denoted by $\langle X_{p},Y_{p}\rangle_{g(p)}$, and the norm of the vector $X_{p}$ computed with the metric $g$ is denoted by $|X_{p}|_{g(p)}=\langle X_{p},X_{p}\rangle_{g(p)}^{1/2}$. Similarly, the gradient of a function $\phi\in C^{\infty}(M)$ is the vector field ${\rm grad}_{g}\,\phi=(\mathbf{d}\phi)^{\sharp_{g}}\in\mathfrak{X}(M)$, where $\mathbf{d}\phi\in\Omega^{1}(M)$ is the exterior derivative of $\phi$ and $\sharp_{g}:T^{*}M\to TM$ denotes the sharp isomorphism, i.e., $\mathbf{d}\phi(X)=\langle{\rm grad}_{g}\,\phi,X\rangle_{g}$ for each $X\in\mathfrak{X}(M)$. In coordinates, it is given by:
@@ 2.1. Configuration space
Let $\ell,n\in\mathbb{N}$. A symmetric, positive-definite matrix $W\in\mathcal{M}_{(n+1)\times(n+1)}(\mathbb{R})$ defines an inner product (and, thus, a Riemannian metric) on $\mathbb{R}^{n+1}$, namely:
$\langle X_{\rm x},Y_{\rm x}\rangle_{g({\rm x})}=X_{\rm x}^{\top}\,W\,Y_{\rm x},\qquad X_{\rm x},Y_{\rm x}\in T_{\rm x}\mathbb{R}^{n+1},~{}{\rm x}\in\mathbb{R}^{n+1},$
where the superscript $\top$ denotes the transpose. The corresponding norm is denoted by $|X_{\rm x}|_{W}=(X_{\rm x}^{\top}\,W\,X_{\rm x})^{1/2}$. The points of $\mathbb{R}^{n+1}$ of unit norm define an $n$-dimensional ellipsoid, which is denoted by:
@@
In this work, we consider a transformer consisting of $\ell$ tokens of dimension $n+1$ constrained to evolve on an ellipsoid. This choice of state space models the effect of token normalization which constrains the “size” of a token as discussed in more detail in the next section. As we have $\ell$ tokens, the resulting state space is the Cartesian product of $\ell$ copies of the ellipsoid, i.e.:
$(\mathcal{E}_{W}^{n})^{\ell}=\underbrace{\mathcal{E}_{W}^{n}\times{\ldots}\times\mathcal{E}_{W}^{n}}_{\ell\text{-times}}.$
Note that $(\mathcal{E}_{W}^{n})^{\ell}$ is an embedded submanifold of:
@@
where $\mathbb{R}_{0}^{n+1}=\mathbb{R}^{n+1}-\{0\}$. The natural inclusion is denoted by $\imath_{W}:(\mathcal{E}_{W}^{n})^{\ell}\hookrightarrow(\mathbb{R}_{0}^{n+1})^{\ell}$ and we define the projection:
@@
where:
(2) $\pi_{W}:\mathbb{R}_{0}^{n+1}\to\mathcal{E}_{W}^{n},\quad{\rm x}\mapsto\pi_{W}({\rm x})={\rm x}\,|{\rm x}|_{W}^{-1}.$
The corresponding tangent map is readily seen to be:
$T_{\mathbf{x}}\boldsymbol{\pi}_{W}=T_{{\rm x}_{1}}\pi_{W}\times\ldots\times T_{{\rm x}_{\ell}}\pi_{W}:T_{\mathbf{x}}(\mathbb{R}_{0}^{n+1})^{\ell}\to T_{\boldsymbol{\pi}_{W}(\mathbf{x})}(\mathcal{E}_{W}^{n})^{\ell},$
for each $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\in(\mathbb{R}_{0}^{n+1})^{\ell}$, where the tangent map of $\pi_{W}$ at each ${\rm x}\in\mathbb{R}_{0}^{n+1}$ is given by:
(3) $T_{\rm x}\pi_{W}:T_{\rm x}\mathbb{R}_{0}^{n+1}\to T_{\pi_{W}({\rm x})}\mathcal{E}_{W}^{n},\quad X_{\rm x}\mapsto T_{\rm x}\pi_{W}\cdot X_{\rm x}=|{\rm x}|_{W}^{-1}\left(\mathbb{I}_{n+1}-{\rm x}\,{\rm x}^{\top}\,W\,|{\rm x}|_{W}^{-2}\right)\cdot X_{{\rm x}}.$
In particular, for ${\rm y}\in\mathcal{E}_{W}^{n}$, we have $T_{\rm y}\pi_{W}\cdot X_{\rm y}=\left(\mathbb{I}_{n+1}-{\rm y}\,{\rm y}^{\top}\,W\right)\cdot X_{\rm y}$.
@@
For each ${\rm x}\in\mathbb{R}^{n+1}$, we make the identification $T_{\rm x}\mathbb{R}^{n+1}\simeq\mathbb{R}^{n+1}$. In particular, for ${\rm y}\in\mathcal{E}_{W}^{n}$, we have:
$T_{\rm y}\mathcal{E}_{W}^{n}=\{Y_{\rm y}\in T_{\rm y}\mathbb{R}^{n+1}\simeq\mathbb{R}^{n+1}\mid{\rm y}^{\top}\,W\,Y_{\rm y}=0\}.$
@@
=$=\left\{Y_{\mathbf{y}}=(Y_{{\rm y}_{1}},\ldots,Y_{{\rm y}_{\ell}})\in(\mathbb{R}^{n+1})^{\ell}\mid{\rm y}_{i}^{\top}\,W\,Y_{{\rm y}_{i}}=0,~{}1\leq i\leq\ell\right\}.$
@@
There are a number of models in which the tokens evolve on the $n$-sphere, $\mathbb{S}^{n}=\mathcal{E}_{\mathbb{I}_{n+1}}^{n}$. For brevity, in that case we will drop the subscripts standing for the matrix $W=\mathbb{I}_{n+1}$. For instance, we will write $|\cdot|=|\cdot|_{\mathbb{I}_{n+1}}$, $\boldsymbol{\pi}=\boldsymbol{\pi}_{\mathbb{I}_{n+1}}$, etc.

### v2

@@ Abstract.
The transformer architecture has become the foundation of modern Large Language Models (LLMs), yet its theoretical properties are still not well understood. As with classic neural networks, a common approach to improve these models is to increase their size and depth. However, such strategies may be suboptimal, as several works have shown that adding more layers yields increasingly diminishing returns. More importantly, prior studies have shown that increasing depth may lead to model collapse, i.e., all the tokens converge to a single cluster, undermining the ability of LLMs to generate diverse outputs. Building on differential equation models for the transformer dynamics, we prove that all the tokens in a transformer asymptotically converge to a cluster as depth increases. At the technical level we leverage tools from control theory, including consensus dynamics on manifolds and input-to-state stability (ISS). We then extend our analysis to auto-regressive models, exploiting their structure to further generalize the theoretical guarantees.
@@ 1. Introduction
The transformer architecture [1], introduced in 2017, was a groundbreaking work that established the foundation for current Large Language Models (LLMs). In contrast to older recurrent architectures [2; 3], transformers were primarily based on the attention mechanism [4], making them significantly more efficient and powerful. Since then, many works have strived to improve the architecture through modifications, such as rotary embeddings [5], although the core structure has remained intact over the years.
Nearly a decade later, this architecture is still not fully understood [6]. Numerous works have attempted to bridge the gap between experimental results and theoretical guarantees, some focusing on the approximation capabilities of transformers [7], others on their in-context behavior [8]. Beyond simply analyzing or corroborating empirical results, many theoretical efforts have aimed to justify or guide the evolution of transformers by studying individual components such as the attention block [9].
Although several theoretical results suggest the key to better transformers may not just be depth [10], the trend toward ever-larger models continues. In fact, recent commercial models, such as the LLaMA 3 405B, have surpassed one hundred layers [11]. This increase in depth, along with the significant number of parameters, has made recent LLMs prohibitively expensive to train and operate [12]. As the cost of transformers continues to rise, this trend underscores a pressing question: does increasing the number of layers actually improve performance?
As more works attempt to answer this question, one conclusion becomes increasingly clear: deeper is not always better. In fact, some empirical studies on recent models suggest that the final layers do not significantly contribute to the model’s representational capacity [13; 14]. These studies corroborate theoretical results showing that the expressive power of transformers increases with depth only up to a certain threshold [10]. Furthermore, as highlighted by [15], training transformers beyond a few hundred layers is often difficult unless architectural modifications are added.
Other works reached a more dooming conclusion: as depth increases, tokens begin to cluster. One of the first studies on the topic [9], proved that pure attention loses rank with depth. In such event the model may collapse, i.e., lose the ability to generate diverse outputs, as the depth increases. However, a common limitation of these works typically lies in the strong assumptions that limit their practical implications. Thus, some studies have used more general models, considering for instance the norm in each layer [16].
Our work focuses on expanding these results by relaxing some of their assumptions, leveraging well-established results from control theory, e.g., consensus [17; 18] and input-to-state stability (ISS) [19]. We build on the mathematical model derived in [20], which describes the transformer dynamics by a differential equation, and prove that all tokens converge to a cluster. By interpreting tokens as particles on the sphere, we leverage the extensive literature on consensus dynamics defined over such manifolds [21]. Additionally, using ISS theory, we extend the results of [20; 22] to auto-regressive models such as GPT-2, while allowing the weight matrices to vary with depth, unlike previous works. Finally, our theoretical predictions are confirmed through experiments on two popular models: GPT-2 and GPT-Neo.
@@
Table 1. Summary of the results presented in this work for several particular cases of the continuous model Eq.˜6, where $Q(t)$, $K(t)$, and $U(t)$ denote the query, key and value matrices, respectively.
Full attention
Section Section˜3 Section˜4 # of heads $h=1$ $h\geq 1$ $P(t)=Q(t)^{\top}K(t)$ Time invariant, symmetric, positive definite Time varying, unif. continuous, bounded $U(t)$ Identity Identity Result Theorem˜3.2 Theorem˜4.1 Statement Gradient flow, convergence to equilibrium Convergence to consensus Domain of attraction Whole sphere Some hemisphere
Causal attention (auto-regressive)
Section Section˜5.1 Section˜5.2 Section˜5.3 # of heads $h\geq 1$ $h=1$ $h=1$ $P(t)=Q(t)^{\top}K(t)$ Time varying, bounded Time varying, bounded Time varying, bounded $U(t)$ Identity Time invariant, symmetric Time varying, bounded, symmetric Result Theorem˜5.1 Theorem˜5.2 Theorem˜5.3 Statement Asympt. stability of consensus Asympt. stability of consensus Convergence to ball around consensus Domain of attraction Conull (compl. of zero measure) Fixed hemisphere Time-varying hemisphere
Contributions of the paper
The main contribution of this paper is to provide a number of results for a differential equation model of attention [^1] Since we focus on the attention mechanism, this model does not describe the effect of feed-forward layers. showing that all tokens converge to a single cluster thereby leading to a collapse of the model. We use the term consensus equilibria to refer to such clusters as is done in the consensus literature [23; 24]. These results hold under different assumptions on the parameters of the model —namely, the query ($Q$), key ($K$) and value matrices ($U$), as well as the number of heads (hh)— that are summarized in Table˜1.
In particular, with Theorem˜3.2 we prove that the dynamics of the transformer is a Riemannian gradient vector field, from which we conclude convergence to an equilibrium point (guaranteed to be of consensus type when $P=Q^{\top}K$ is the identity) for every initial position of the tokens. Although the gradient nature of the dynamics, in this case, was already observed and exploited in [25], for the benefit of the readers we provide a formal proof of this fact in a slightly more general setting. Theorem˜4.1 states that tokens converge to a consensus equilibrium whenever their starting positions lie in the interior of some hemisphere of the ellipsoid. This result holds for any number of heads and time varying matrix $P=Q^{\top}K$ provided that $U$ is the identity and $P$ is bounded and uniformly continuous as a function of the time. A similar result is reported in [25] under Lemma 4.2. However, its conclusions hold under the stronger assumptions that both $U$ and $P=Q^{\top}K$ are the identity matrix and there is a single attention head.
Theorems˜3.2 and 4.1 make no assumptions on the attention matrix other those induced by the assumptions on $P$. In contrast, Theorems˜5.1, 5.2 and 5.3 focus on the auto-regressive case, also known as causal attention, where the self-attention matrix is lower triangular. Theorem˜5.1 states that when $U$ is the identity, the first token is fixed and all the other tokens converge to the position of the first one for almost every initial position of the tokens. In fact, we have asymptotic stability of this consensus equilibrium. This holds for any number of heads and any time varying $P$ matrix provided it is bounded. Similar conclusions are reported under Theorem 4.1 in [26] by imposing stronger assumptions: time invariance of $P=Q^{\top}K$ and existence of a single attention head. Theorem˜5.2 extends these result to the case where $U$ is a time invariant symmetric matrix and the multiplicity of its largest eigenvalue is one. In this case all the tokens will converge to a consensus equilibrium (moreover, that equilibrium is asymptotically stable) if they start in one of the two hemispheres defined by the eigenvector associated with the largest eigenvalue of $U$. We were only able to establish this result for the single-head case although we believe it holds in greater generality. To the best of the authors’ knowledge there is no result available in the literature for the case where $U$ is not the identity matrix although this is conjectured, but not proved, in [26]. Lastly, in Theorem˜5.3 we consider the case where $U(t)$ is time varying. Thus, the eigenvector corresponding to the largest eigenvalue is also time varying, and the result states that the tokens will converge to a ball around the consensus whose radius depends on how fast this eigenvector moves. The convergence takes place provided all tokens start on the hemisphere determined by this eigenvector at the initial time. This paper extends the authors’ previous work in [27] by incorporating two additional scenarios: the gradient flow and the time-varying value matrix. In addition, it provides complete proofs of all results.
Our theoretical findings are validated using experiments with the GPT-2 and the GPT-Neo models, providing empirical evidence for convergence to consensus equilibria in more general situations than those captured by our theoretical results. Thus, demonstrating additional confirmation for model collapse.
Notations and conventions
@@
We use the letters $n,\ell,r,$ and $s$ to denote elements of $\mathbb{N}=\{1,2,\ldots\}$. The space of $r\times s$ real matrices is denoted by $\mathcal{M}_{r\times s}(\mathbb{R})$. In particular, $\mathbb{I}_{r}\in\mathcal{M}_{r\times r}(\mathbb{R})$ denotes the identity matrix. The Frobenius norm of a square matrix $A\in\mathcal{M}_{r\times r}(\mathbb{R})$ is denoted by $\|A\|$. The elements of $\mathbb{R}^{n+1}$ are regarded as column matrices, i.e., ${\rm x}\in\mathbb{R}^{n+1}\equiv\mathcal{M}_{(n+1)\times 1}(\mathbb{R})$. When it is convenient, tuples of $\ell$ elements of $\mathbb{R}^{n+1}$, $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\in(\mathbb{R}^{n+1})^{\ell}$ (note the different font), will be regarded either as matrices, $\mathbf{x}\in\mathcal{M}_{(n+1)\times\ell}(\mathbb{R})$, or column matrices, $\mathbf{x}\in\mathcal{M}_{(n+1)\ell\times 1}(\mathbb{R})$. The tangent space of a smooth manifold $M$ at $p\in M$ and its elements are denoted by $T_{p}M$ and $X_{p}\in T_{p}M$, respectively. The tangent bundle and the space of vector fields of a smooth manifold $M$ are denoted by $\pi_{M}:TM\to M$ and $\mathfrak{X}(M)=\Gamma(\pi_{M})$, respectively. In the same vein, the space of $k$-forms is denoted by $\Omega^{k}(M)$. Let $M$ be a smooth manifold. The inner product between the vectors $X_{p}\in T_{p}M$ and $Y_{p}\in T_{p}M$, according to a Riemannian metric $g$ on $M$, is denoted by $\langle X_{p},Y_{p}\rangle_{g(p)}$, and the norm of the vector $X_{p}$ computed with the metric $g$ is denoted by $|X_{p}|_{g(p)}=\langle X_{p},X_{p}\rangle_{g(p)}^{1/2}$. Similarly, the gradient of a function $\phi\in C^{\infty}(M)$ is the vector field ${\rm grad}_{g}\phi=(\mathbf{d}\phi)^{\sharp_{g}}\in\mathfrak{X}(M)$, where $\mathbf{d}\phi\in\Omega^{1}(M)$ is the exterior derivative of $\phi$ and $\sharp_{g}:T^{*}M\to TM$ denotes the sharp isomorphism, i.e., $\mathbf{d}\phi(X)=\langle{\rm grad}_{g}\phi,X\rangle_{g}$ for each $X\in\mathfrak{X}(M)$. In coordinates, it is given by:
@@
Given another smooth manifold $N$, the tangent map of a smooth map $\phi\in C^{\infty}(M,N)$ is denoted by $T\phi:TM\to TN$ while its pullback is denoted by $\phi^{*}:T^{*}N\to T^{*}M$.
@@ 2.1. Configuration space
Let $\ell,n\in\mathbb{N}$. A symmetric, positive-definite matrix $W\in\mathcal{M}_{(n+1)\times(n+1)}(\mathbb{R})$ defines an inner product on $\mathbb{R}^{n+1}$:
$\langle X_{\rm x},Y_{\rm x}\rangle_{W}=X_{\rm x}^{\top}WY_{\rm x},$
for each $X_{\rm x},Y_{\rm x}\in T_{\rm x}\mathbb{R}^{n+1}$ and ${\rm x}\in\mathbb{R}^{n+1}$, where the superscript $\top$ denotes the transpose. The corresponding norm is denoted by $|X_{\rm x}|_{W}=(X_{\rm x}^{\top}WX_{\rm x})^{1/2}$. The points of $\mathbb{R}^{n+1}$ of unit norm define an $n$-dimensional ellipsoid, which is denoted by:
@@
In this work, we consider a transformer consisting of $\ell$ tokens of dimension $n+1$ constrained to evolve on an ellipsoid. As we have $\ell$ tokens, the resulting state space is the Cartesian product of $\ell$ copies of the ellipsoid, i.e.:
$(\mathcal{E}_{W}^{n})^{\ell}=\underbrace{\mathcal{E}_{W}^{n}\times{\ldots}\times\mathcal{E}_{W}^{n}}_{\ell\text{-times}},$
which is an embedded submanifold of:
@@
where $\mathbb{R}_{0}^{n+1}=\mathbb{R}^{n+1}-\{0\}$. The natural inclusion is denoted by $\imath_{W}:(\mathcal{E}_{W}^{n})^{\ell}\hookrightarrow(\mathbb{R}_{0}^{n+1})^{\ell}$ and we define the projection as:
@@
with $\pi_{W}:\mathbb{R}_{0}^{n+1}\to\mathcal{E}_{W}^{n}$ given by $\pi_{W}({\rm x})={\rm x}|{\rm x}|_{W}^{-1}$ for each ${\rm x}\in\mathbb{R}_{0}^{n+1}$. The corresponding tangent map is readily seen to be:
$T_{\mathbf{x}}\boldsymbol{\pi}_{W}:T_{\mathbf{x}}(\mathbb{R}_{0}^{n+1})^{\ell}\to T_{\boldsymbol{\pi}_{W}(\mathbf{x})}(\mathcal{E}_{W}^{n})^{\ell},\quad T_{\mathbf{x}}\boldsymbol{\pi}_{W}=T_{{\rm x}_{1}}\pi_{W}\times\ldots\times T_{{\rm x}_{\ell}}\pi_{W},$
for each $\mathbf{x}=({\rm x}_{1},\ldots,{\rm x}_{\ell})\in(\mathbb{R}_{0}^{n+1})^{\ell}$, with $T_{\rm x}\pi_{W}:T_{\rm x}\mathbb{R}_{0}^{n+1}\to T_{\pi_{W}({\rm x})}\mathcal{E}_{W}^{n}$ given by:
(2) $T_{\rm x}\pi_{W}\cdot X_{\rm x}=|{\rm x}|_{W}^{-1}\left(\mathbb{I}_{n+1}-{\rm x}{\rm x}^{\top}W|{\rm x}|_{W}^{-2}\right)\cdot X_{{\rm x}},$
for each ${\rm x}\in\mathbb{R}_{0}^{n+1}$ and $X_{\rm x}\in T_{\rm x}\mathbb{R}_{0}^{n+1}$. In particular, for ${\rm y}\in\mathcal{E}_{W}^{n}$ and $X_{\rm y}\in T_{\rm y}\mathcal{E}_{W}^{n}$, we have:
$T_{\rm y}\pi_{W}\cdot X_{\rm y}=\left(\mathbb{I}_{n+1}-{\rm y}{\rm y}^{\top}W\right)\cdot X_{\rm y}.$
@@
For each ${\rm x}\in\mathbb{R}_{0}^{n+1}$, we make the identification $T_{\rm x}\mathbb{R}_{0}^{n+1}\simeq\mathbb{R}^{n+1}$. In particular, for ${\rm y}\in\mathcal{E}_{W}^{n}$, we have:
$T_{\rm y}\mathcal{E}_{W}^{n}=\{Y_{\rm y}\in T_{\rm y}\mathbb{R}_{0}^{n+1}\simeq\mathbb{R}^{n+1}\mid{\rm y}^{\top}WY_{\rm y}=0\}.$
@@
$=\left\{Y_{\mathbf{y}}=(Y_{{\rm y}_{1}},\ldots,Y_{{\rm y}_{\ell}})\in(\mathbb{R}^{n+1})^{\ell}\mid{\rm y}_{i}^{\top}WY_{{\rm y}_{i}}=0,~1\leq i\leq\ell\right\}.$
@@
There are a number of models in which the tokens evolve on the $n$-sphere, i.e., $\mathbb{S}^{n}=\mathcal{E}_{\mathbb{I}_{n+1}}^{n}$. For brevity, in that case we will drop the subscripts standing for the matrix $W=\mathbb{I}_{n+1}$. For instance, we will write $|\cdot|=|\cdot|_{\mathbb{I}_{n+1}}$, $\boldsymbol{\pi}=\boldsymbol{\pi}_{\mathbb{I}_{n+1}}$, etc.